OpenStudy (anonymous):

Marco found five charms he likes. Using your function determine the cost of the bracelet he would make.

OpenStudy (anonymous):

I am not sure how many he would be able to make but my function is f(x)=85+15x

OpenStudy (anonymous):

any answers given

OpenStudy (anonymous):

I really need help

OpenStudy (anonymous):

give me a min

OpenStudy (anonymous):

okay thanx :)

OpenStudy (anonymous):

opposite of add is subtract from both sides and it gives u x=-70

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

where did u get 85

OpenStudy (anonymous):

so x = -70 is how many bracelets he made?

OpenStudy (anonymous):

I got it from this: Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15.

OpenStudy (anonymous):

so after reading that I made the function and now I need to use my function to see how many bracelets he would make.

OpenStudy (anonymous):

2 sisters purchase 2 charms and 1 bracelet =$115 answer 2 charms

OpenStudy (anonymous):

did it help u

OpenStudy (anonymous):

Well yes that could be used as an example but the original question is: Marco found five charms he likes. Using your function determine the cost of the bracelet he would make. So we figured out how many two would cost, so how many would five cost?

OpenStudy (anonymous):

15*5=$75 75+85=$160 5 charms is $75

OpenStudy (anonymous):

Okay that helps a lot, thank you! can I ask one more question?

OpenStudy (anonymous):

yes just fan me and medal

OpenStudy (anonymous):

How do I do that? I am new to openstudy. :P

OpenStudy (anonymous):

Nvm I did it.

OpenStudy (anonymous):

cool thanks thats it helping someone else

OpenStudy (anonymous):

?

OpenStudy (anonymous):

No i am good I think i can figure the question out myself. thank for the help

OpenStudy (anonymous):

no problem see u later